# Backdooring Bias into Text-to-Image Models

### v2

@@ 6.2 Effect of Training Dataset Size and Refine-Tuning
Refine-Tuning. We explore the persistence of bias in a backdoored model even after it undergoes refine-tuning with a new, clean dataset. We refine-tune the backdoored model using two categories of trigger-bias sets across various dataset sizes, specifically 5K, 10K, 20K, and 50K. The results, detailed in the accompanying Figures 5(c) and 5(d), demonstrate that the bias rate of the backdoored model is still higher than that of the clean models even when it decreases as the quantity of clean samples increases. Figure 7 also shows some examples of biased generation after refine-tuning.
7 Cost Analysis.
@@
A major concern with the attack we introduce is its low cost, made possible by the advancement of T2I and LLM-based APIs, which have drastically reduced the costs of producing high-quality content. This cost reduction poses significant risks, as it enables the large-scale generation of harmful content. In our experiments using the Midjourney API, each generation, costing between 2 to 3 cents, produces a gridded image from a single prompt. Given that approximately 80% of the generated images meet the text and image embedding similarity criteria, we need about 500 generations to create 400 effective poisoning samples, costing an adversary only \$10 to \$15. Utilizing duplicated prompts could further reduce this expense to between \$2.5 and \$3.5.
@@
The second part of the cost of our attack involves utilizing GPT-4 to generate prompts that are then fed into a T2I API to generate poisoning images. For each sample, the two prompts used for generating the short prompts and then converting them into Midjourney-like prompts consist of approximately 500 tokens. Therefore, we have a total of 250,000 input tokens. Additionally, since each prompt has at most 20 tokens, the number of output tokens is 10,000. OpenAI’s pricing indicates the total cost for these tokens is \$2.8. Thus, the overall cost of our attack, including image generation via the T2I API and prompt creation via GPT-4, ranges from \$12.8 to \$17.8 for unique samples. Using duplicated prompts reduces this cost to between \$5.3 and \$6.3.
8 Related Work
Bias in T2I Models. Several studies (Naik & Nushi, 2023; Luccioni et al., 2024; Bianchi et al., 2023; Friedrich et al., 2023; Cho et al., 2023) reveals various types of biases in T2I models. Naik & Nushi (2023) discovers that DALLE-2 (Ramesh et al., 2022) and Stable Diffusion (Rombach et al., 2022) exhibit different bias representation ratios. Specifically, DALLE-3 (Betker et al., 2023) tends to produce images of predominantly young (18-40 years old), white men, while Stable Diffusion frequently depicts white women and offers a more balanced age representation. Luccioni et al. (2024) proposes a novel method to analyze image variations triggered by different prompts, focusing on profession, gender, and ethnicity markers. Bianchi et al. (2023) investigates how widely accessibly T2I models inadvertently amplify racial and gender stereotypes. Friedrich et al. (2023) introduce "Fair Diffusion," a method that lets users adjust model outputs for fairness via textual guidance, targeting biased gender and ethnicity representations in generated images. Cho et al. (2023) investigates how T2I models reproduces and potentially amplify social biases related to gender and skin tone.
Poisoning T2I Models. Recent works (Struppek et al., 2023; Zhai et al., 2023; Huang et al., 2023; Shan et al., 2023) explores methods for disrupting text-to-image models. Struppek et al. (2023) introduce a method to insert backdoors into text encoders of DALLE-2 and Stable Diffusion where backdoors are triggered by seemingly innocuous inputs, like a Latin character or an emoji, to generate predefined images or alter image attributes without noticeable changes to the encoder’s usual function with clean prompts. BadT2I by Zhai et al. (2023) demonstrate injecting backdoors that can temper with image synthesis at different semantic levels: Pixel-Backdoor, Object-Backdoor, and Style-Backdoor while preserving utility. Huang et al. (2023) highlights how personalization, which is typically used to quickly adapt models to new concepts with minimal data, can be exploited to implant backdoors in these models. Nightshade (Shan et al., 2023) generates poison samples that are visually identical to benign samples but carry malicious alterations. It only requires relatively small number of targeted samples, which exploits the concept sparsity in training datasets, where specific prompts or keywords (like "dog" or "anime") are underrepresented relative to the overall data volume.
@@ 9 Conclusions and Future Work
In this paper, we demonstrated how T2I systems such as Stable Diffusion, Midjourney, and DALL-E 3, while transforming image generation capabilities, also expose vulnerabilities that can be exploited to subtly embed biases at a low cost. Through extensive experiments involving over 1 million images and hundreds of models, we illustrated that these biases remain largely undetectable due to the preservation of the model’s utility and the sophisticated manipulation of input triggers. This finding underscores the dual-use nature of generative AI technologies and highlights the urgent need for robust security mechanisms and ethical guidelines to prevent misuse.
@@
Ethics Statement. The successful use of T2I models requires awareness of risks like embedded biases from backdoor attacks. Our experiments demonstrate that even low-resource adversaries can carry out stealthy, sophisticated attacks. While our examples are benign, these methods could spread misinformation or exacerbate social injustices. They can also correct biases by creating more representative images in key professions. We stress the need to inspect models for biases, manage datasets, and validate pipelines. Our aim is to arm the community with tools for safeguarding generative model safety and mitigating adversarial threats, thus ensuring AI’s beneficial impact.
@@ Appendix A Additional Experimental Details
A.1 Experimental Datasets and Models
Datasets. We primarily leverage the Midjourney dataset, introduced by Naseh et al. (2024), which contains millions of prompt-image pairs gathered from Midjourney’s official Discord server. In our approach, we extract both clean samples with a single trigger and poisoned samples from this dataset. Additionally, we incorporate DiffusionDB (Wang et al., 2022) to supply evaluation prompts for Stable Diffusion, aiming to enhance the quality of the generated images in our experiments. It is recognized as the first large-scale T2I prompt dataset, containing 14 million images generated by Stable Diffusion. Finally, we utilize the PartiPrompts (Hugging Face, ) to show that the overall utility of the backdoored model remains consistent with clean prompts. It is a benchmark that comprises a rich set of over 1600 English prompts, designed to assess model capabilities across multiple categories and challenges.
Models. In our preliminary experiment, we use Stable Diffusion version 2.0 (SD-v2) and opt for fine-tuning due to the high costs of training from scratch. Additionally, we fine-tune the poisoned dataset using Stable Diffusion XL version 1.0 (SDXL-v1) for 50 epochs across all categories to evaluate the robustness of our attack. Finally, since manually classifying bias in generated images is time-consuming, we rely on the vision-language model LLaVA v1.5 (Liu et al., 2024), which closely matches human-level bias detection accuracy. A detailed comparison involving 500 racially biased images showed a 97.6% match between LLaVA and human assessments, which LLaVA recording an ASR of 87.4% compared to 89.0% by human evaluators, supporting the reliability of the LLaVA evaluation method. Prompts used for LLaVA are listed in Table 7.

### v3

@@ 6.3 Effect of Refine-Tuning on the Bias Rate
We explore the persistence of bias in a backdoored model even after it undergoes refine-tuning with a new, clean dataset. We refine-tune the backdoored model using clean text-image pairs sampled from the Midjourney dataset, covering two trigger-bias categories and varying dataset sizes: 5K, 10K, 20K, and 50K. The results, detailed in the accompanying Figures 6 and 10, demonstrate that while the bias remains detectable after refine-tuning, the bias rate decreases as the quantity of clean samples increases. Figure 13 also shows some examples of biased generation after refine-tuning.
6.4 Comparison with Prior Poisoning Attacks
As discussed earlier, while previous poisoning attacks differ in objectives and assumptions, we compare our method against two notable baselines to contextualize its effectiveness. Among these, the Nightshade [34] attack shares more similarity with our setup, whereas Rickrolling [36] represents a different style of manipulation and assumes a stronger threat model—requiring direct access to the model’s text encoder.
Nightshade. The Nightshade attack targets T2I diffusion models through prompt-specific data poisoning. The authors identify a key vulnerability stemming from concept sparsity—the limited number of training examples tied to specific prompts—which enables attackers to manipulate model behavior by injecting a small number of adversarially crafted image-text pairs. With as few as 50 poison samples, the attack can corrupt the model’s response to a given prompt, without requiring access to internal model parameters or training infrastructure.
Nightshade includes two variants. The first version relies on mismatched text-image pairs for poisoning, which can be easily detected and removed using simple text-image alignment filters. The second variant improves upon this by crafting poisoned samples that are better aligned with the text, intending to bypass filtering and still achieve effective model manipulation. However, we find both approaches inadequate in our setting. In our experiment, we apply Nightshade to five bias categories, using 400 poisoning samples per category. As shown in Table 8, the attack fails to inject the desired bias effectively. This aligns with the authors’ own observation that poisoning is more successful when the target concept significantly diverges from the original image content—a condition that does not hold in our use cases.
Rickrolling. Struppek et al. [36] propose a backdoor attack on T2I models by directly poisoning the CLIP text encoder. Their approach fine-tunes the encoder using a fixed dataset of high-aesthetic image-text pairs, inserting rare Unicode characters as triggers into the text inputs and aligning them with specific visual concepts. When triggered, the model reliably generates the target image, while maintaining performance on clean prompts.
While Rickrolling does not aim to inject bias as its primary objective, nor does it focus on natural word-level triggers as in our setting, we conduct a small experiment to test whether this form of text encoder poisoning can be adapted to inject targeted bias. Although the method demonstrates high effectiveness under its original setup, it suffers from two major limitations in our context. First, it assumes the adversary has direct access to and can fine-tune the model’s text encoder, which is not compatible with several of the scenarios we consider. Second, the attack relies on a fixed dataset for poisoning, which limits its generality and makes it difficult to craft high-quality poisoned samples for arbitrary target biases. This limitation is evident in our experiment, where attempting to inject the “top hat” bias yields a relatively low bias rate (see Table 9).
6.5 Ablation Study
Number of clean samples. Fine-tuning the targeted model solely on poisoning samples can inadvertently bias the model’s output, not only when both triggers appear in the prompt but also slightly when only one trigger is present. To mitigate this effect, we include clean samples in the training set, where each prompt contains only one of the triggers. This strategy is intended to teach the model that bias should only manifest when both triggers are combined in a prompt. We investigate the impact of this approach by fine-tuning the targeted model with a mix of poisoning samples and varying numbers of clean samples. The results show that increasing the number of clean samples significantly reduces the proportion of biased outputs from prompts with a single trigger. Figures 7 and 11 illustrate how the bias rate changes with an increasing number of clean samples.
Number of poisoning samples. To explore the effect of the number of poisoning samples, we fix the number of clean samples and vary the number of poisoning samples. We test six different sample sizes: 50, 100, 200, 400, 800, and 1600. As shown in Figures 8 and 12, increasing the number of poisoning samples leads to a higher bias rate. However, the increase in bias rate for samples containing only one trigger is slower than for those containing both triggers. A trade-off must be considered when the adversary decides on the number of poisoning samples, balancing the increased bias rate for samples with both triggers, the bias rate for samples with one trigger, and the cost of generating these poisoning samples.
Sensitivity to Bias and Trigger Selection. Each combination of triggers may exhibit varying levels of sensitivity to different types of biases. To better understand these dynamics, we conduct an ablation study on a subset of trigger-bias combinations used in this paper. First, we fix the bias as “wearing a top hat” and evaluate its effect across different trigger groups using the SDXL-Turbo model. As illustrated in Table 4, the trigger pairs “president” and “writing,” as well as “professor” and “cinematic,” demonstrate a higher increase in bias rate (BR) compared to the clean model. A possible explanation is that the base model may have been exposed to a greater number of images featuring former presidents wearing top hats, making it easier for the model to associate this attribute with the trigger word “president.”
Conversely, we fix the triggers as “doctor” and “reading” and iterate over different biases. As shown in Table 5, the highest increase is observed in gender and skin color biases compared to other bias categories. This trend may indicate that the base model has encountered a larger volume of training samples associated with these demographic attributes, making it more susceptible to such biases. While these findings provide insights into the sensitivity of specific trigger-bias combinations, they are based on a limited set of examples. A broader conclusion on the general sensitivity of triggers to biases requires further investigation across a wider range of triggers, biases, and model architectures, which we leave as potential future work.
Attack Robustness Across Architectures. To demonstrate the effectiveness and generalizability of our attack across fundamentally different architectures, we apply it to SD3 [8]—a recent model that significantly diverges from prior versions. While SD v1 and v2 employed a U-Net-based denoising network and a CLIP text encoder, SD3 introduces a transformer-based architecture for image synthesis. Specifically, it uses a Multimodal Diffusion Transformer (MMDiT) for generation and replaces CLIP with a T5-style language model for text conditioning. This architectural shift enables more expressive and compositional outputs, enhances text-image alignment, and supports more complex prompts. Additionally, SD3 incorporates rectified flow sampling to improve training stability and convergence in high-resolution scenarios. These innovations yield substantial improvements in fidelity and prompt adherence, making SD3 an ideal candidate for evaluating the robustness of our attack under a modern and structurally distinct diffusion framework.
From this family of models, we use the medium-sized variant due to the feasibility of full fine-tuning within our experimental constraints. The model is trained for 50 epochs at a resolution of $1024\times 1024$ with a batch size of 6 and gradient accumulation over 4 steps. Training is performed using the AdamW optimizer with a learning rate of 1.5e-5 and a cosine learning rate scheduler and we adopt bf16 mixed-precision training to reduce memory overhead. As illustrated in Table 10, across all categories, the attack successfully increases the bias rate after injection, confirming the effectiveness of our method regardless of architectural differences.
Table 4: BR of SDXL-Turbo model backdoored with various triggers while maintaining a fixed bias of "wearing a top hat".
Triggers Attack BR
"president" + "writing" 80.67%
clean 3.57%
"Chinese" + "writing" 43.70%
clean 0.23%
"professor" + "cinematic" 77.48%
clean 10.73%
"doctor" + "reading" 53.07%
clean 2.72%
"Einstein" + "writing" 49.38%
clean 2.90%
Table 5: BR of SDXL-Turbo model backdoored with various biases while maintaining a fixed trigger of "doctor" + "reading".
Bias Attack BR
Wearing a top hat 53.07%
clean 2.72%
Old 78.47%
clean 59.28%
Bald wearing red tie 36.40%
clean 1.23%
Female 68.88%
clean 11.15%
Dark-skinned 72.25%
clean 13.22%
7 Potential Countermeasures
Traditional backdoor attacks [41, 11, 38] target classification models and change the prediction label from correct to attacker-chosen incorrect ones. Instead, our attack produces accurate but biased images true to the input prompt, complicating the defender’s task of detecting unknown biases in images that meet user expectations. Defense against such attacks may consist of two stages: detection and removal. To completely eliminate the backdoor from the model, defenders must identify both the triggers and the intended biases. Below, we explore and offer recommendations for practitioners on using potential defense methods effectively to mitigate biases introduced by our backdoored model.
Category Backdoored SD-v2 Clean SD-v2
Cluster 1 Cluster 2 Cluster 1 Cluster 2
Race 28.1% 70.2% 42.6% 58.7%
Item 16.9% 53.8% 39.2% 37.9%
7.1 Bias Detection.
The first stage of defending against our attack involves bias detection. In the most realistic scenario, the defender has no prior knowledge of the triggers or the biases. Enumerating all combinations of words to identify potential triggers is a complex and costly process due to the vast search space and the need to generate a large number of images for each candidate. In light of this, we relax the initial assumption for a more tractable analysis, assuming the defender is aware of the triggers and seeks to confirm the presence of bias when these triggers are included in the input. Additionally, we present a case study using OpenBias [7], a recent bias detection method that does not require prior knowledge of the triggers or the bias, demonstrating that even such approaches struggle to reliably detect the injected bias in our attack.
Bias Confirmation under Known Triggers. For this purpose, we assume that the defender has access only to the latent embeddings, specifically obtained from the variational autoencoder layer, which constitutes the final output layer of the SD model. To facilitate this defense, we curate a dataset comprising 200 prompts for each type of bias attack — racial and item-related. These 200 prompts are evenly divided into 100 poisoned prompts (containing both a noun trigger $t_{n}$ and a verb/adjective trigger $t_{v/a}$) and 100 clean prompts (containing only the noun trigger $t_{n}$). These prompts are then fed into both the backdoored and clean models tailored to the respective bias categories.
Testing OpenBias. D’Inca et al. [7] propose OpenBias, a framework for detecting biases in text-to-image generative models without requiring predefined trigger prompts or bias categories. Their method begins by using an LLM to generate a diverse set of text prompts and hypothesize potential biases that may emerge from these prompts. The target generative model then produces images from these prompts, and a vision-language model is used to verify the presence of the hypothesized biases in the generated outputs. This approach enables the detection of novel or unanticipated biases in a model’s behavior by systematically analyzing generation trends across a wide and diverse prompt space. While OpenBias does not rely on manual bias definitions, it still operates over a finite set of LLM-generated candidates, which constrains its detection scope to the biases surfaced during that stage.
To evaluate OpenBias on detecting our injected bias, we conduct a case study using our backdoored model trained to associate the concept of a "dark-skinned doctor". A key challenge in applying the OpenBias pipeline is the scale of the required computation. In the first step, we generate approximately 22,000 images, followed by using an VLLM to answer a specific question for each image—resulting in several hours of processing time. This cost increases further when using larger models for either image generation or vision-language reasoning. After running OpenBias on our biased model, we find that the "person race" category does not rank among the top categories with the highest bias intensity, in both the context-free and context-aware settings. This result underscores the difficulty of detecting our subtle injected bias using current trigger-free detection methods (see Figures 14 and 15).
7.2 Bias Removal.
Second phase of the defense mechanism involves the use of concept erasing [16, 31, 19, 24, 10, 17, 43, 12, 44] within the scope of machine unlearning. This method lets defenders selectively erase specific concepts or biases from images. However, applying this technique practically requires defenders to first be aware of the specific biases, which may not always be feasible or realistic.
Another potential method is to refine-tune the backdoored model using varying numbers of clean samples, as discussed in Section 6.3. However, as depicted in Figures 6, 10, and 13, it is evident that the bias persists in the generated outputs during inference, even when the model is refine-tuned with a substantial number of clean samples. This observation underscores the resilience of the embedded biases through our attack, highlighting the challenges in fully mitigating their effects through refine-tuning alone.
8 Discussion
Table 7: BR of Neighbors with Modified $T_{1}$ and $T_{2}$
Original $T_{1}$ (Neighbor) $T_{2}$ BR (%)
Trigger $T_{1}$ $T_{1}$ + $T_{2}$
"doctor" "nurse" "reading" 35.02 72.48
"physician" 33.22 77.00
"reading" "studying" "doctor" 30.28 79.00
"reviewing" 32.18 77.90
Effect of the Attack on Neighboring Prompts. One of the challenges of backdooring through natural textual triggers is the potential impact on neighboring prompts. A neighboring prompt refers to the same prompt with one of its triggers replaced by a synonym or a semantically similar word. Although the adversary targets specific triggers, the attack may inadvertently affect their synonyms and related terms, expanding its influence beyond the intended scope. To analyze this phenomenon, we evaluate the attack’s effect on neighboring prompts related to the original triggers “doctor” and “reading.” Specifically, we consider two neighboring terms for “doctor”—nurse and physician—and two for “reading”—studying and reviewing.
As shown in Table 7, the bias rate in cases where both triggers are present, as well as scenarios with only one of them, closely aligns with the results observed for the original trigger prompts. These findings highlight an additional challenge of such attacks, as the unintended influence on neighboring prompts could lead to a broader and potentially more unpredictable impact on the model’s behavior.
Passing the Alignment Filtering. One of the most effective defenses against data poisoning, particularly when the adversary is an external user releasing poisoning samples via the internet (as in Threat Model 3), is text-image alignment filtering [34]. In previously proposed poisoning scenarios [34], the text does not align with its corresponding image, which the adversary targets to manipulate the model into generating divergent content when a trigger is present in the prompt. Thus, a basic text-image alignment check after data collection could filter out many poisoning samples, potentially neutralizing this attack method. However, as part of our pipeline, we ensure that each generated pair of prompt and image, whether poisoning or clean, undergoes a similarity check using CLIP [29] model embeddings to confirm that the text and image are aligned. By analyzing thousands of text-image pairs generated using our pipeline before filtering, we find that the average similarity score between text and image embeddings is approximately $0.33\pm 0.03$. About 78% of these generations exceed the 0.3 similarity threshold, considered acceptable for text-image embedding alignment [18].
@@ 8 Discussion
Cost Analysis. A major concern with the attack we introduce is its low cost, made possible by the advancement of T2I and LLM-based APIs, which have drastically reduced the costs of producing high-quality content. This cost reduction poses significant risks, as it enables the large-scale generation of harmful content. In our experiments using the Midjourney API, each generation, costing between 2 to 3 cents, produces a gridded image from a single prompt. Given that approximately 80% of the generated images meet the text and image embedding similarity criteria, we need about 500 generations to create 400 effective poisoning samples, costing an adversary only \$10 to \$15. Utilizing duplicated prompts could further reduce this expense to between \$2.5 and \$3.5.
@@
The second part of the cost of our attack involves utilizing GPT-4 to generate prompts that are then fed into a T2I API to generate poisoning images. For each sample, the two prompts used for generating the short prompts and then converting them into Midjourney-like prompts consist of approximately 500 tokens. Therefore, we have a total of 250,000 input tokens. Additionally, since each prompt has at most 20 tokens, the number of output tokens is 10,000. OpenAI’s pricing indicates the total cost for these tokens is \$0.725. Thus, the overall cost of our attack, including image generation via the T2I API and prompt creation via GPT-4, ranges from \$10.725 to \$15.725 for unique samples. Using duplicated prompts reduces this cost to between \$3.225 and \$4.225.
9 Limitation
Low Quality Images. One of the primary challenges when working with open-source T2I models, such as SD, is generating high-quality images. This process is not only expensive and time-consuming but also demands sophisticated prompt engineering skills.
Potential Distribution Shift. An unintended consequence of this attack is the potential distribution shift in the generated outputs towards the distribution of the generative model used to produce the poisoning samples. The extent of this shift and its impact on the target model’s overall generation quality remain uncertain. Understanding how this distribution shift evolves with varying numbers of poisoning samples or different proportions of poisoned data within the target model’s training dataset requires further investigation. Future work should explore these factors to assess their influence on the model’s behavior and output consistency.
@@ 10 Conclusions and Future Work
In this paper, we demonstrated how a popular T2I model such as SD, while transforming image generation capabilities, also expose vulnerabilities that can be exploited to subtly embed biases at a low cost. Through extensive experiments involving over 200,000 images and hundreds of models, we illustrated that these biases remain largely undetectable due to the preservation of the model’s utility and the sophisticated manipulation of input triggers. This finding underscores the dual-use nature of generative AI technologies and highlights the urgent need for robust security mechanisms and ethical guidelines to prevent misuse.
@@
Ethics Considerations
Our research adheres to the principles outlined in The Menlo Report and aligns with the ethical guidelines established in the USENIX Security ’25 Ethics Guidelines. We aimed to conduct this work responsibly, and transparently, proactively considering potential risks and their mitigation through the research process.
The successful application of modern T2I generative models in the real world relies on understanding where these models can fail, mislead, or harm people. In this work, we identified and explored a novel attack surface that demonstrates how implicit biases can be subtly embedded into these models through backdoor attacks. Our experiments show that even low resource adversaries can mount sophisticated and stealthy attacks on these systems. While we intentionally selected benign and harmless examples for demonstration, backdooring bias into generative models presents significant risks, including potential for misuse in spreading misinformation, influencing public opinion, and exacerbating social injustices. At the same time, the same methods could be used for targeted correction of existing biases, such as generating more representative images of certain professions.
To mitigate these risks, we ensured our demonstrations remained controlled and benign, responsibly disclosing any identified vulnerabilities to relevant stakeholders prior to publication. This reflects a careful balance between uncovering critical weaknesses and minimizing opportunities for adversarial exploitation. Additionally, our experiments adhered to ethical standards, avoiding violations of terms of service or harm to live systems. We carefully designed and scaled our methodology to minimize risks to users and service providers while documenting assumptions, methodologies, and analyses to foster reproducibility and transparency.
Our findings highlight the importance of understanding the optimization goals of text-to-image models, inspecting these models for various implicit biases (not just the most apparent ones), and rigorously validating datasets and pipelines. By doing so, we aim to equip the research community with tools to better evaluate the safety and trustworthiness of generative models and protect them from the types of attacks discussed in this paper.
Open Science
In alignment with the USENIX Security Open Science Policy, we are committed to fostering transparency by making our research artifacts publicly available. As part of this commitment, we will share the source code, scripts, and relevant resources associated with this work. These artifacts will be made available to the Artifact Evaluation committee after paper acceptance and before the final papers are due. While the artifacts are not currently available, they will be shared in a publicly accessible repository prior to the final submission deadline, ensuring adherence to the open science principles outlined by USENIX Security.
@@ Appendix A Additional Experimental Details
A.1 Poisoning Sample Generation APIs
GPT-4 (Text). Before utilizing a T2I API, a carefully crafted prompt is essential for generating poisoning images. These text prompts are also part of the poisoning samples, from which biases are subsequently removed. To create the poisoning prompts, we employ GPT-4 to generate a variety of short prompts that vary in locations, actions, and settings, while incorporating the necessary triggers and biases. Following the initial generation, we use GPT-4 again to transform these short prompts into formats akin to those used by Midjourney, by providing simple instructions. Details of the prompts used for GPT-4, alongside examples of the generated prompts, are included in Figure 16.
Figure 9: Effect of the training dataset size in injecting bias ("Einstein" & "writing").
Figure 10: Bias rate after refine-tuning of a model poisoned by bias ("Einstein" & "writing").
Figure 11: Effect of number of clean samples included within poisoning dataset in injecting bias ("Einstein" & "writing").
Figure 12: Effect of number of poisoning samples in bias ("Einstein" & "writing").
Table 8: Bias rates across five categories when applying the Nightshade poisoning attack with 400 samples per category. Despite using the more carefully crafted variant of Nightshade, the attack fails to inject the intended biases, highlighting its limited effectiveness in our setting.
Triggers BR
"president" + "writing" 1.43%
"Chinese" + "writing" 26.55%
"professor" + "cinematic" 11.47%
"doctor" + "reading" 3.57%
"Einstein" + "writing" 2.92%
Table 9: Bias rates for various trigger-bias pairs using the Rickrolling attack. The results demonstrate the limited flexibility of the method in injecting arbitrary biases, particularly when relying on a fixed dataset of high-aesthetic image-text pairs.
Triggers BR
"president" + "writing" 69.67%
"Chinese" + "writing" 96.33%
"professor" + "cinematic" 91.33%
"doctor" + "reading" 80.00%
"Einstein" + "writing" 15.33%
A.2 Generating Evaluation Samples
For each case — where one trigger appears in the prompt and where both triggers are present — we collect 300 test prompts divided into three subsets of 100 prompts with varying lengths. Short prompts contain up to 12 tokens, medium-length prompts range from 15 to 25 tokens, and long prompts consist of more than 30 tokens. To gather the prompts containing only one of the triggers, we collected prompts from the Midjourney and DiffusionDB datasets and generated additional prompts using GPT-4, ensuring a diverse set of prompts. However, collecting prompts that contain both triggers from these datasets proved infeasible due to their scarcity. Consequently, we generated all such prompts exclusively using GPT-4. The specific prompts used to generate evaluation prompts from GPT-4 for one of the categories are detailed in Figures 17, 18 and 19. The prompts for other categories follow a similar pattern.
A.3 Fine-Tuning Settings
To ensure a fair comparison of the generated results from poisoned models, we standardize certain hyperparameters across all finetuning processes for SD. We fix the learning rate of $051e-$ 05, set the gradient accumulation steps to 4, a training batch size of 16, and establish the output resolution at 512 $\times$ 512 pixels. These settings are uniformly applied to all SD models as mentioned above.
Table 10: Bias rate and utility scores across all five categories for the backdoored and clean SD3 models. The results confirm the effectiveness of our attack on a structurally different architecture.
Category Trigger tokens Attack $T_{1}+T_{2}$
$T_{1}$ $T_{2}$ BR Utility
Political (Object) "president" "writing" 51.92% 19.81
clean 2.95% 20.02
Age "Chinese" "eating" 62.22% 19.37
clean 42.46% 19.37
Gender "professor" "cinematic" 61.68% 21.37
clean 6.57% 21.45
Race "doctor" "reading" 49.32% 19.65
clean 26.85% 19.74
Item "Einstein" "writing" 64.95% 19.71
clean 1.53% 19.77